How many mL of 0.1M HCl is required to react completely with 1g of mixture of Na2CO3 and NaHCO3 containing equi-molar amounts of both ?

Grade:11

2 Answers

varun
6 years ago
HELLO RAVI,
Let the amount of Na2CO3 in the mixture will be gram
Then,amount of NaHCO3 in the mixture will be (1-x) gram
molar mass of Na2CO3 = 106g/mol
therefore,
number of moles of Na2CO3 = x/106mol
molar mass of NaHCO3 = 84 g/mol
therefore,
number of moles of NaHCO3 = 1-x/84 mol
ACCORDING TO THE QUESTION,
 
X/106 = 1-X/54
84x = 106 – 106x
190x = 106
x = 0.5579
therefore,
number of moles of Na2CO3 = 0.5579/106 mol
                                              = 0.0053mol
And number of moles of NaHCO3 = 1 – 0.5579/84
                                                    = 0.0053 mol
Hcl reacts with Na2CO3 and Na2CO3 according to the following equation
 
 
2HCl + Na2CO3 ------------> 2NaCl + H2O + CO2
 2mol      1mol                  
and
HCl + NaHCO3 ----------->NaCl  + H2O + CO2
1mol     1 mol
so,
1 mol of Na2CO3 reacts with 2 mol of HCl 
therefore,0.0053 mol of Na2CO3 reacts with 2X0.0053 mol = 0.0106mol
and,
1 mol of NaHCO3 reacts with 1 mol of HCl 
therefore,0.0053 mol of NaHCO3 reacts with 0.0053 mol of HCl
total moles of HCl = 0.0106 + 0.0053
                           =  0.0159 moles
In 0.1 mol HCL 
0.1 mol of HCl is present in 1000ml of solution
therefore,
0.0159 mol of HCL is present in 
1000 X 0.0159/0.1
 = 159 ml of solution
Hence 159 ml of HCl is required to react completely with 1g mixture of Na2CO3 and NaHCO3 containing equimolar amounts of both
THANK YOU
 
Ajeet Tiwari
3 years ago
hello student

Answer:

158 mL

Explanation:

let the amount of Na₂CO₃ be x

& that of NaHCO₃ be 1-x

Now moles of Na₂CO₃ = x / 106

& moles of NaHCO₃ = 1-x / 84

Now according to question , number of moles of Na₂Co₃ = number of moles of NaHCO₃

Therefore x / 106 = 1-x / 84

84x = 106-106x

84x +106x = 106

190x = 106

Or

x = 106 / 190 = 0.558

Therefore moles of Na₂Co₃ = 0.558 / 106 = 0.00526

& moles of NaHCO₃ = 1 - 0.558 / 84 = 0.0053

Now Hcl reacts with Na₂Co₃ & NaHCO₃ as follows:



From the above reactions, 1 mol of Na₂Co₃ will react with 2 mol of Hcl

Therefore 0.00526 mol of Na₂Co₃ will react with 2 x 0.00526 mol of Hcl & similarly 0.00526 mol of NaHCO₃ will react with 0.00526 mol of Hcl

Total moles of Hcl required to react with mixture of of NaHCO₃ & Na₂Co₃

= 2 X 0.00526 + 0.00526 =0.01578 mol

Also according to question 0.1 mol of 0.1 M Hcl is present in 1000 ml

Or

0.01578 mol of 0.1 M Hcl is present in (1000/0.1) x 0.01578 = 158 ml

Hence, 158 mL of 0.1 M of HCl is required to react completely with 1 g mixture of Na₂Co₃ and NaHCO₃, containing equimolar amounts of both.

hope it helps
thankyou